Find the x-intercept of the graph of y=3+6x
Answer by longjonsilver(2297)

the x-intercept... where the line crosses the x-axis. The x-axis is the line y=0, so put this piece of information in. We get:

y=3+6x
0=3+6x
6x = - 3
x = -3/6
x = -1/2

jon.